OpenStudy (anonymous):

Find the quotient: four-ninths divided by eight. one-eighteenth eighteen three and five-ninths nine over thirty-two

OpenStudy (texaschic101):

(4/9) / 8 = 4/9 * 1/8 = 4/72 = 1/18 The 8 is the same as 8/1 and when dividing fractions, you flip what you are dividing by and then multiply.